Opened 10 years ago

Closed 9 years ago

Last modified 2 years ago

#259 closed flight processing (fixed)

EU09/06, flight day 222/2009, Vatna

Reported by: knpa Owned by: knpa
Priority: alpha 4 medium Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by mark1)

Data location: ~arsf/arsf_data/2009/flight_data/ipy/EU09_06-2009_222_Vatna

Data arrived from ARSF via SATA disk Daphne03 on 24/8/09

Scientific objective: Recording glacial responses to climatic forcing

Priority: Alpha 4-Medium

PI: D. Graham

Sensors:

  • Eagle (13/05/2010)
  • Leica LIDAR (13/05/2010)
  • Digital Photography (6/04/2010)

Change History (18)

comment:1 Changed 9 years ago by mark1

photograph delivery created and ready for checking.

comment:2 Changed 9 years ago by mark1

  • Description modified (diff)

comment:3 Changed 9 years ago by mark1

  • Description modified (diff)

photography dispatched to PI on 6/04/2010

comment:4 Changed 9 years ago by mark1

  • Owner set to mark1
  • Status changed from new to accepted

Started Hyperspectral processing

comment:5 Changed 9 years ago by mark1

Paused hyperspectral processing and started LIDAR processing - need a DEM to process the hyperspectral.

comment:6 Changed 9 years ago by mark1

Eagle lines 2 and 9 have no specim sync.

comment:7 Changed 9 years ago by mark1

LIDAR and hyperspectral delivery directories created and ready for checking at:

~airborne/workspace/EU09_06-2009_222_Vatna/delivery/20100505/EU09_06
~airborne/workspace/EU09_06-2009_222_Vatna/delivery/20100507/EU09_06

comment:8 Changed 9 years ago by knpa

Delivery checked lidar (.../delivery/20100505/EU09_06)

All looks good apart from following:

  • Readme needs to be unix2dos'd
  • dem/README_DEM_MAY_NOT_BE_ACCURATE.txt needs to be unix2dos'd
  • can see two flightlines with major problems (cloud/ice?) and two with minor - should mention in readme ?)
  • min/max time given as 0 from check_ascii_lidar.sh for flightlines: 3-17.

comment:9 Changed 9 years ago by mark1

Above corrections have been applied and ascii files re-created to fix the time problem.

comment:10 Changed 9 years ago by knpa

Delivery checked hyperspectral (.../delivery/20100507/EU09_06)

All fine apart from:

  • No misc folder - intentional?
  • no dem - is the one in lidar delivery appropriate to use or is there a SRTM patched version we can include?
  • No bin listing in Readme contents
  • FastQC claims there are no dark frames...
  • Need to unix2dox Read_Me
  • e13 has some gaps in lev3 screenshot - mention in Readme?

comment:11 Changed 9 years ago by mark1

Made the relevant above changes to hyperspectral delivery.

comment:12 Changed 9 years ago by mark1

Eagle SCTs

line sct
1 0.0
2 -3.01
3 0.0
4 0.0
5 0.0
6 0.0
7 0.0
8 -0.03
9 -1.65
10 -0.06
11 0.0
12 0.0
13 0.0
14 0.0
15 0.0
16 0.0
17 0.0

comment:13 Changed 9 years ago by mark1

  • Description modified (diff)

Data dispatched to PI on 13/05/2010

comment:14 Changed 9 years ago by mark1

  • Component changed from Processing: general to Archiving

rsynced data over to the repository (3/6/2010).

comment:15 Changed 9 years ago by knpa

  • Owner changed from mark1 to knpa
  • Status changed from accepted to assigned

Transferred to, and backed up by, NEODC.

Moved from repository to archive disk.

comment:16 Changed 9 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Removing workspace copy and closing ticket.

comment:17 Changed 2 years ago by lah

Archiving

Tidied and ready for upload to NEODC.

comment:18 Changed 2 years ago by asm

Archiving

Re-processed hyperspectral data has been archived and everythin is ready to be downloaded from CEDA. Will close this ticket.

Note: See TracTickets for help on using tickets.